OpenStudy (anonymous):

what is 1000 plus 5480 divided by 2 squared?

OpenStudy (anonymous):

1000+5480 = 6480 2 squared = 4 6480/4 = 1620

OpenStudy (anonymous):

depends on which you perform first. this could mean [(1000+5480)/2]^2 =10,497,600 or [(1000+5480)/2^2]=1620